OpenStudy (anonymous):

assume f(x) = -2x + 8 and g(x)=3x what is the value of (f+g)(-1)

OpenStudy (anonymous):

f+g= (-2x+8)+(3x)=x+8 so (f+g)(-1)=(-1)+8=7

OpenStudy (anonymous):

I thought it was (8)(-1) so basically -8 ?

OpenStudy (anonymous):

I mean -2x+8+3x =0 x= -8 -8(-1) = 8

OpenStudy (anonymous):

oh now you don't set equal to 0 you just add f and g together -2x+8+3x x+8 which gives you the equation that you plug -1 into (-1)+8

OpenStudy (anonymous):

Thank youu , can you help with another question too ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

assume f(x) = 2x + 8 and g(x) = 5 what is the value of (gof) (3)